Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for oneHuron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below: Standard Quantity or Hours Standard Price or Rate Standard Cost Direct materials 6.00 pounds $ 2.30 per pound $ 13.80 Direct labor 0.50 hours $ 10.00 per hour $ 5.00 During the most recent month, the following activity was recorded: Eleven thousand pounds of material were purchased at a cost of $2.10 per pound. All of the material purchased was used to produce 1,500 units of Zoom. 500 hours of direct labor time were recorded at a total labor cost of $6,500. Required: 1. Compute the materials price and quantity variances for the month. 2. Compute the labor rate and efficiency variances for the month.
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (2.30 - 2.10)*11,000

= 2200 F

Materials Quantity Variance = (SQ-AQ) *SP

= (1500*6 - 11000)*2.30

= 4600 U

Labor Rate Variance = (SR-AR) *AH

= (10 - 6500/500)*500

= 1500 U

Labor efficiency variance = (SH-AH) *SR

= (1500*0.50 - 500)*10

= 2500 F

COMMENT IF YOU FACE ANY ISSUES